# HyPINO: Multi-Physics Neural Operators via HyperPINNs and the Method of Manufactured Solutions

### v1

@@ Abstract
We present HyPINO, a multi-physics neural operator designed for zero-shot generalization across a broad class of parametric PDEs without requiring task-specific fine-tuning. Our approach combines a Swin Transformer-based hypernetwork with mixed supervision: (i) labeled data from analytical solutions generated via the Method of Manufactured Solutions (MMS), and (ii) unlabeled samples optimized using physics-informed objectives. The model maps PDE parametrizations to target Physics-Informed Neural Networks (PINNs) and can handle linear elliptic, hyperbolic, and parabolic equations in two dimensions with varying source terms, geometries, and mixed Dirichlet/Neumann boundary conditions, including interior boundaries. HyPINO achieves strong zero-shot accuracy on seven benchmark problems from PINN literature, outperforming U-Nets, Poseidon, and Physics-Informed Neural Operators (PINO). Further, we introduce an iterative refinement procedure that compares the physics of the generated PINN to the requested PDE and uses the discrepancy to generate a "delta" PINN. Summing their contributions and repeating this process forms an ensemble whose combined solution progressively reduces the error on six benchmarks and achieves over 100x gain in average $L_{2}$ loss in the best case, while retaining forward-only inference. Additionally, we evaluate the fine-tuning behavior of PINNs initialized by HyPINO and show that they converge faster and to lower final error than both randomly initialized and Reptile-meta-learned PINNs on five benchmarks, performing on par on the remaining two. Our results highlight the potential of this scalable approach as a foundation for extending neural operators toward solving increasingly complex, nonlinear, and high-dimensional PDE problems with significantly improved accuracy and reduced computational cost.
@@ 1 Introduction
Neural operators have emerged as a promising paradigm for solving partial differential equations (PDEs). Their ability to generalize across families of PDEs, fast inference, and full differentiability make them appealing for a wide range of scientific computing tasks. In the longer term, such methods may serve as building blocks for general-purpose, foundational, and multi-physics simulators, sometimes referred to as "world model predictors" Ding et al. (2022); Li et al. (2025); Wu et al. (2024); Yuan et al. (2025).
@@
However, existing neural operators are typically sample inefficient Herde et al. (2024). As a result, most prior work focuses on narrowly defined problem families Lu et al. (2021a). Variations are limited to singular aspects such as specific PDE parameters (e.g., diffusion coefficients) Cho et al. (2024), boundary conditions Duvall et al. (2021), or domain shapes Zheng et al. (2024). The support for simultaneous variations of PDE operators remains limited to subdomains, such as parametrized convection-diffusion-reaction PDEs Ye et al. (2024).
@@
One way to address the data requirement is by incorporating physics-informed losses. Such losses can provide self-supervision without requiring labeled simulation data Raissi et al. (2019); Wang et al. (2021c). While promising, existing methods often suffer from spectral bias Wang et al. (2022) and mode collapse Yang and Perdikaris (2018). Moreover, purely physics-based training is unstable in practice Goswami et al. (2023). Therefore, even with physics-based losses, obtaining a large, labelled dataset that spans a diverse range of PDEs remains a significant bottleneck.
@@
To overcome these challenges, we propose HyPINO. A hybrid framework that combines physics-informed learning with synthetic supervised data. Our approach leverages a Swin Transformer-based hypernetwork de Avila Belbute-Peres et al. (2021); Ha et al. (2016); Liu et al. (2021) to map PDE specifications to the parameters of a target physics-informed neural network (PINN). Our HyPINO enables zero-shot generalization without task-specific fine-tuning. A key contribution of our work is a scalable synthetic data pipeline that generates two complementary types of training data: (i) Supervised samples generated via the Method of Manufactured Solutions (MMS) Oberkampf et al. (1995) by selecting target solutions and deriving the corresponding PDEs analytically. These provide direct supervision with known reference solutions. (ii) Physics-only samples constructed by randomly sampling PDE operators, source terms, and boundary / initial conditions. These are trained using physics-informed losses without requiring ground-truth solutions. This hybrid training strategy allows us to cover a broad spectrum of two-dimensional linear elliptic, hyperbolic, and parabolic PDEs that cover a wide range of phenomena in the natural and engineering sciences, including heat diffusion, wave propagation, acoustic scattering, and membrane deformation. Furthermore, HyPINO supports problems with mixed Dirichlet and Neumann boundary conditions, as well as complex domain geometries. In addition, we introduce an iterative refinement procedure that builds an ensemble of corrective PINNs. At each iteration, the model evaluates the residual error and fits a "delta" PINN to improve the solution. This ensemble refinement provides a lightweight alternative to traditional fine-tuning, requiring only inference passes rather than full backward passes.
@@
We evaluate HyPINO on seven diverse PDE benchmarks, demonstrating improved zero-shot generalization compared to baselines such as U-Nets Ronneberger et al. (2015), Poseidon Herde et al. (2024), and PINO Li et al. (2024). We also find that PINNs initialized with our method fine-tune more efficiently than those starting from random or meta-learned initializations, achieving faster convergence and lower final errors.
@@ 2 Related Work
Neural operators aim to approximate solution operators that map PDE specifications to continuous solution fields, enabling fast, mesh-free inference and generalization to unseen problem instances Hao et al. (2023a); Li et al. (2022, 2020, 2024); Lu et al. (2021a); Marwah et al. (2023). Recent work scales these ideas toward foundation models that ingest large corpora of simulated data or equation specifications and promise broad cross-task transfer Hao et al. (2024); Herde et al. (2024); McCabe et al. (2023); Subramanian et al. (2024); Ye et al. (2024). Despite rapid progress, most operators still target narrow PDE families (e.g. fixed equations with varying coefficients) and depend on expensive high-fidelity solvers for supervision Ye et al. (2024). Embedding the governing equations in the loss function alleviates the need for labeled data and improves physical fidelity Raissi et al. (2019); Wang et al. (2021c). While the original formulation was introduced for stand-alone Physics-Informed Neural Networks (PINNs), the same residual losses have recently been integrated into operator architectures, yielding Physics-Informed Neural Operators (PINO) that train from unlabeled residual samples Boudec et al. (2024); Goswami et al. (2023); Li et al. (2024). These approaches still require careful weighting of supervision terms and often struggle with stability and spectral bias for complex PDEs.
Hypernetworks generate the parameters of a target network conditioned on an auxiliary input Ha et al. (2016). In the PDE context, HyperPINNs predict PINN weights for varying coefficients de Avila Belbute-Peres et al. (2021); Lee et al. (2023), and subsequent works extend this idea to boundary conditions, domain changes, and low-rank weight modulation Cho et al. (2023); Duvall et al. (2021); Hadorn (2022); Morel et al. (2025). Yet, existing models rarely support concurrent variation of multiple operators, geometries, and boundary types without task-specific fine-tuning.
The Method of Manufactured Solutions (MMS) provides analytic ground-truth pairs by choosing a target field and deriving the corresponding source term and boundary data Oberkampf et al. (1995). MMS has long served for numerical-solver verification and was recently adopted for PINN evaluation Kashefi and Mukerji (2022) and operator training Hasani and Ward (2024). However, prior studies focus on single equations (e.g. Poisson); leveraging MMS for multi-physics operator pre-training remains largely unexplored.
@@ 3.1 PDE Parameterization
To support a wide range of linear PDEs while maintaining compatibility with modern machine learning models, we adopt a parameterization that is flexible, user-friendly, and efficiently processed by state-of-the-art architectures. The function $f$ is discretized on a uniform grid over $\Omega$, resulting in a 2D array $F$ representing its values at grid points. The boundary conditions are parametrized by creating two 2D grids per boundary type ($\partial\Omega_{D}$, $\partial\Omega_{N}$): (i) A binary mask $M$ indicating the presence of the boundary at each grid point, where we assign a value of 1 to the four grid cells closest to each boundary point and zero elsewhere; and (ii) a value grid $V$ storing the corresponding boundary values ($g$ for Dirichlet or $h$ for Neumann conditions) at those marked cells, with zeros elsewhere. Figure 2 illustrates a sampled PDE instance with its full parameterization. Finally, following Iwata et al. (2023), we parameterize $\mathcal{L}$ as $\mathcal{L}[u](\mathbf{x})=c_{1}u+c_{2}u_{x}+c_{3}u_{y}+c_{4}u_{xx}+c_{5}u_{yy}$, where $\mathbf{c}=(c_{1},c_{2},c_{3},c_{4},c_{5})\in\mathbb{R}^{5}$ encodes the operator coefficients.
@@ 3.2 Neural Operator Architecture
We base our model on HyperPINN de Avila Belbute-Peres et al. (2021), a hypernetwork-based neural operator that maps a parametrized PDE instance to the weights $\theta$ of a PINN $u_{\theta}$ specialized to that instance. Formally, the hypernetwork realizes a mapping
@@
The vector of operator coefficients $\mathbf{c}\in\mathbb{R}^{5}$ is embedded into a fixed-length representation $z_{C}\in\mathbb{R}^{d_{C}}$ using a Fourier feature mapping Tancik et al. (2020) which was shown to prevent spectral bias and mode collapse, in particular in physics-informed settings Wang et al. (2021b). The grid-based inputs $F$, $M_{g}$, $M_{h}$, $V_{g}$ and $V_{h}$ are concatenated and processed via a series of $K$ Swin Transformer blocks $\{\mathcal{SW}_{i}\}_{i=1}^{K}$ Liu et al. (2021). After each block, we introduce a FiLM layer Perez et al. (2017), which modulates the Swin block’s output conditioned on where $z_{C}$:
@@
Inspired by Swin Transformer UNet architectures Cao et al. (2021); Fan et al. (2022), we retain all intermediate latent representations $\{z_{i}\}_{i=1}^{K}$ to keep information at various semantic levels. To enable information aggregation, we flatten the spatial dimensions $H_{i}$ and $W_{i}$ and use Multi-Head Attention Pooling Lee et al. (2019); Zhai et al. (2022), where a set of trainable query vectors $\{q_{i}\}_{i=1}^{K},q_{i}\in\mathbb{R}^{T\times C_{i}}$ is defined. $T$ corresponds to the number of weight and bias tensors in the target PINN. The queries $q_{i}$ are then used in a multi-head attention module together with $z_{i}$ reshaped into $kv_{i}\in\mathbb{R}^{H_{i}\times W_{i}\times C_{i}}$:
@@
We define the architecture of the target PINN as MLP with Fourier feature mapping Tancik et al. (2020), which, when concatenated to the input $(x,y)$, results in a dimensionality of $2N+2$, and multiplicative skip connections Wang et al. (2021a). Fourier encodings provide spectral expressivity for modeling high-frequency components Wang et al. (2021b), while the skip connections enhance gradient propagation and, in the context of hypernetworks, have the additional benefit of enabling dynamic depth modulation based on PDE complexity by allowing the hypernet to mask some layers.
@@ 3.3 Data Sampling
Domains $\Omega\subset[-1,1]^{2}$ are generated via randomized Constructive Solid Geometry (CSG) Lu et al. (2021c). The outer boundary $\partial\Omega_{\text{outer}}$ is defined as the unit square and may represent either spatial or spatio-temporal domains, with y=1y=-1 corresponding to the initial time in time-dependent PDEs. Inner boundaries $\partial\Omega_{\text{inner},i}$ are formed by subtracting randomly sampled geometric primitives (e.g., disks, polygons, rectangles) from the outer region. These boundaries enclose regions where the source term $f$ remains active, and their role (e.g., obstacle vs. inclusion) is encoded implicitly through the boundary type.
@@ 3.4 Objective Function
is the residual loss and $\rho(\cdot)$ the Huber function Huber (1992). The Dirichlet and Neumann losses are computed similarly:
@@
For PDEs with known analytical solutions $u$, we add a second-order Sobolev loss Czarnecki et al. (2017) that penalizes errors in function values, gradients, and second derivatives:
@@ 3.5 Residual–Driven Iterative Refinement
Using a hypernetwork to generate a single PINN of fixed architecture may seem restrictive, particularly in multi-physics settings where different PDEs may demand different levels of representational complexity. However, hypernetworks offer a natural mechanism for generating ensembles of PINNs at inference time, which have proven effective in reducing prediction error Bischof and Kraus (2022); Jagtap and Karniadakis (2020); Stiller et al. (2020). Beyond naïvely producing multiple independent samples, our framework for linear PDEs supports an ensemble construction through an iterative refinement procedure, similar in spirit to multi-stage neural networks that progressively reduce residual error Wang and Lai (2024):
@@ 4.2 Baseline Models
We compare HyPINO against three baselines, each trained for 30,000 batches with batch size 128 and an initial learning rate of $10^{-4}$: (i) UNet Ronneberger et al. (2015), which shares HyPINO’s encoder but replaces the hypernetwork decoder with a convolutional decoder that directly outputs a $224\times 224$ solution grid. It is trained solely on supervised data and has 62M trainable parameters. (ii) Poseidon Herde et al. (2024), a pretrained neural operator with 158M parameters. We use the Poseidon-B checkpoint and adapt it by changing the embedding and lead-time-conditioned layer normalization layers’ dimensionality to match the size of our parametrization. Similarly to the UNet, Poseidon is fine-tuned only on supervised data. (iii) PINO Li et al. (2024), a Fourier neural operator Li et al. (2020) with 33M parameters. We adapt it to accept 5-channel grid inputs and condition on the PDE operator using FiLM layers. It is trained using the same hybrid supervision and curriculum as HyPINO, including physics-informed losses.

### v3

@@ Abstract
We present HyPINO, a multi-physics neural operator designed for zero-shot generalization across a broad class of parametric PDEs without requiring task-specific fine-tuning. Our approach combines a Swin Transformer-based hypernetwork with mixed supervision: (i) labeled data from analytical solutions generated via the Method of Manufactured Solutions (MMS), and (ii) unlabeled samples optimized using physics-informed objectives. The model maps PDE parametrizations to target Physics-Informed Neural Networks (PINNs) and can handle linear elliptic, hyperbolic, and parabolic equations in two dimensions with varying source terms, geometries, and mixed Dirichlet/Neumann boundary conditions, including interior boundaries. HyPINO achieves strong zero-shot accuracy on seven benchmark problems from PINN literature, outperforming U-Nets, Poseidon, and Physics-Informed Neural Operators (PINO). Further, we introduce an iterative refinement procedure that compares the physics of the generated PINN to the requested PDE and uses the discrepancy to generate a "delta" PINN. Summing their contributions and repeating this process forms an ensemble whose combined solution progressively reduces the error on six benchmarks and achieves over 100x gain in average $L_{2}$ loss in the best case, while retaining forward-only inference. Additionally, we evaluate the fine-tuning behavior of PINNs initialized by HyPINO and show that they converge faster and to lower final error than both randomly initialized and Reptile-meta-learned PINNs on five benchmarks, performing on par on the remaining two. Our results highlight the potential of this scalable approach as a foundation for extending neural operators toward solving increasingly complex, nonlinear, and high-dimensional PDE problems. The code and model weights are publicly available at https://github.com/rbischof/hypino.
@@ 1 Introduction
Neural operators have emerged as a promising paradigm for solving partial differential equations (PDEs). Their ability to generalize across families of PDEs, fast inference, and full differentiability make them appealing for a wide range of scientific computing tasks. In the longer term, such methods may serve as building blocks for general-purpose, foundational, and multi-physics simulators, sometimes referred to as "world model predictors" [9, 26, 51, 54].
@@
However, existing neural operators are typically sample inefficient [19]. As a result, most prior work focuses on narrowly defined problem families [31]. Variations are limited to singular aspects such as specific PDE parameters (e.g., diffusion coefficients) [6], boundary conditions [10], or domain shapes [56]. The support for simultaneous variations of PDE operators remains limited to subdomains, such as parametrized convection-diffusion-reaction PDEs [53].
@@
One way to address the data requirement is by incorporating physics-informed losses. Such losses can provide self-supervision without requiring labeled simulation data [41, 48]. While promising, existing methods often suffer from spectral bias [49] and mode collapse [52]. Moreover, purely physics-based training is unstable in practice [12]. Therefore, even with physics-based losses, obtaining a large, labelled dataset that spans a diverse range of PDEs remains a significant bottleneck.
@@
To overcome these challenges, we propose HyPINO. A hybrid framework that combines physics-informed learning with synthetic supervised data. Our approach leverages a Swin Transformer-based hypernetwork [8, 13, 30] to map PDE specifications to the parameters of a target physics-informed neural network (PINN). Our HyPINO enables zero-shot generalization without task-specific fine-tuning. A key contribution of our work is a scalable synthetic data pipeline that generates two complementary types of training data: (i) Supervised samples generated via the Method of Manufactured Solutions (MMS) [39] by selecting target solutions and deriving the corresponding PDEs analytically. These provide direct supervision with known reference solutions. (ii) Physics-only samples constructed by randomly sampling PDE operators, source terms, and boundary / initial conditions. These are trained using physics-informed losses without requiring ground-truth solutions. This hybrid training strategy allows us to cover a broad spectrum of two-dimensional linear elliptic, hyperbolic, and parabolic PDEs that cover a wide range of phenomena in the natural and engineering sciences, including heat diffusion, wave propagation, acoustic scattering, and membrane deformation. Furthermore, HyPINO supports problems with mixed Dirichlet and Neumann boundary conditions, as well as complex domain geometries. In addition, we introduce an iterative refinement procedure that builds an ensemble of corrective PINNs. At each iteration, the model evaluates the residual error and fits a "delta" PINN to improve the solution. This ensemble refinement provides a lightweight alternative to traditional fine-tuning, requiring only inference passes rather than full backward passes.
@@
We evaluate HyPINO on seven diverse PDE benchmarks, demonstrating improved zero-shot generalization compared to baselines such as U-Nets [42], Poseidon [19], and PINO [29]. We also find that PINNs initialized with our method fine-tune more efficiently than those starting from random or meta-learned initializations, achieving faster convergence and lower final errors.
@@ 2 Related Work
Neural operators aim to approximate solution operators that map PDE specifications to continuous solution fields, enabling fast, mesh-free inference and generalization to unseen problem instances [15, 27, 28, 29, 31, 35]. Recent work scales these ideas toward foundation models that ingest large corpora of simulated data or equation specifications and promise broad cross-task transfer [17, 19, 36, 44, 53]. Despite rapid progress, most operators still target narrow PDE families (e.g. fixed equations with varying coefficients) and depend on expensive high-fidelity solvers for supervision [53]. Embedding the governing equations in the loss function alleviates the need for labeled data and improves physical fidelity [41, 48]. While the original formulation was introduced for stand-alone Physics-Informed Neural Networks (PINNs), the same residual losses have recently been integrated into operator architectures, yielding Physics-Informed Neural Operators (PINO) that train from unlabeled residual samples [3, 12, 29]. These approaches still require careful weighting of supervision terms and often struggle with stability and spectral bias for complex PDEs.
Hypernetworks generate the parameters of a target network conditioned on an auxiliary input [13]. In the PDE context, HyperPINNs predict PINN weights for varying coefficients [8, 24], and subsequent works extend this idea to boundary conditions, domain changes, and low-rank weight modulation [5, 10, 14, 37]. Yet, existing models rarely support concurrent variation of multiple operators, geometries, and boundary types without task-specific fine-tuning.
The Method of Manufactured Solutions (MMS) provides analytic ground-truth pairs by choosing a target field and deriving the corresponding source term and boundary data [39]. MMS has long served for numerical-solver verification and was recently adopted for PINN evaluation [23] and operator training [18]. However, prior studies focus on single equations (e.g. Poisson); leveraging MMS for multi-physics operator pre-training remains largely unexplored.
@@ 3.1 PDE Parameterization
To support a wide range of linear PDEs while maintaining compatibility with modern machine learning models, we adopt a parameterization that is flexible, user-friendly, and efficiently processed by state-of-the-art architectures. The function $f$ is discretized on a uniform grid over $\Omega$, resulting in a 2D array $F$ representing its values at grid points. The boundary conditions are parametrized by creating two 2D grids per boundary type ($\partial\Omega_{D}$, $\partial\Omega_{N}$): (i) A binary mask $M$ indicating the presence of the boundary at each grid point, where we assign a value of 1 to the four grid cells closest to each boundary point and zero elsewhere; and (ii) a value grid $V$ storing the corresponding boundary values ($g$ for Dirichlet or $h$ for Neumann conditions) at those marked cells, with zeros elsewhere. Figure 2 illustrates a sampled PDE instance with its full parameterization. Finally, following [21], we parameterize $\mathcal{L}$ as $\mathcal{L}[u](\mathbf{x})=c_{1}u+c_{2}u_{x}+c_{3}u_{y}+c_{4}u_{xx}+c_{5}u_{yy}$, where $\mathbf{c}=(c_{1},c_{2},c_{3},c_{4},c_{5})\in\mathbb{R}^{5}$ encodes the operator coefficients.
@@ 3.2 Neural Operator Architecture
We base our model on HyperPINN [8], a hypernetwork-based neural operator that maps a parametrized PDE instance to the weights $\theta$ of a PINN $u_{\theta}$ specialized to that instance. Formally, the hypernetwork realizes a mapping
@@
The vector of operator coefficients $\mathbf{c}\in\mathbb{R}^{5}$ is embedded into a fixed-length representation $z_{C}\in\mathbb{R}^{d_{C}}$ using a Fourier feature mapping [45] which was shown to prevent spectral bias and mode collapse, in particular in physics-informed settings [47]. The grid-based inputs $F$, $M_{g}$, $M_{h}$, $V_{g}$ and $V_{h}$ are concatenated and processed via a series of $K$ Swin Transformer blocks $\{\mathcal{SW}_{i}\}_{i=1}^{K}$ [30]. After each block, we introduce a FiLM layer [40], which modulates the Swin block’s output conditioned on where $z_{C}$:
@@
Inspired by Swin Transformer UNet architectures [4, 11], we retain all intermediate latent representations $\{z_{i}\}_{i=1}^{K}$ to keep information at various semantic levels. To enable information aggregation, we flatten the spatial dimensions $H_{i}$ and $W_{i}$ and use Multi-Head Attention Pooling [25, 55], where a set of trainable query vectors $\{q_{i}\}_{i=1}^{K},q_{i}\in\mathbb{R}^{T\times C_{i}}$ is defined. $T$ corresponds to the number of weight and bias tensors in the target PINN. The queries $q_{i}$ are then used in a multi-head attention module together with $z_{i}$ reshaped into $kv_{i}\in\mathbb{R}^{H_{i}\times W_{i}\times C_{i}}$:
@@
We define the architecture of the target PINN as MLP with Fourier feature mapping [45], which, when concatenated to the input $(x,y)$, results in a dimensionality of $2N+2$, and multiplicative skip connections [46]. Fourier encodings provide spectral expressivity for modeling high-frequency components [47], while the skip connections enhance gradient propagation and, in the context of hypernetworks, have the additional benefit of enabling dynamic depth modulation based on PDE complexity by allowing the hypernet to mask some layers.
@@ 3.3 Data Sampling
Domains $\Omega\subset[-1,1]^{2}$ are generated via randomized Constructive Solid Geometry (CSG) [33]. The outer boundary $\partial\Omega_{\text{outer}}$ is defined as the unit square and may represent either spatial or spatio-temporal domains, with y=1y=-1 corresponding to the initial time in time-dependent PDEs. Inner boundaries $\partial\Omega_{\text{inner},i}$ are formed by subtracting randomly sampled geometric primitives (e.g., disks, polygons, rectangles) from the outer region. These boundaries enclose regions where the source term $f$ remains active, and their role (e.g., obstacle vs. inclusion) is encoded implicitly through the boundary type.
@@ 3.4 Objective Function
is the residual loss and $\rho(\cdot)$ the Huber function [20]. The Dirichlet and Neumann losses are computed similarly:
@@
For PDEs with known analytical solutions $u$, we add a second-order Sobolev loss [7] that penalizes errors in function values, gradients, and second derivatives:
@@ 3.5 Residual–Driven Iterative Refinement
Using a hypernetwork to generate a single PINN of fixed architecture may seem restrictive, particularly in multi-physics settings where different PDEs may demand different levels of representational complexity. However, hypernetworks offer a natural mechanism for generating ensembles of PINNs at inference time, which have proven effective in reducing prediction error [1, 22, 43]. Beyond naïvely producing multiple independent samples, our framework for linear PDEs supports an ensemble construction through an iterative refinement procedure, similar in spirit to multi-stage neural networks that progressively reduce residual error [50]:
@@ 4.2 Baseline Models
We compare HyPINO against three baselines, each trained for 30,000 batches with batch size 128 and an initial learning rate of $10^{-4}$: (i) UNet [42], which shares HyPINO’s encoder but replaces the hypernetwork decoder with a convolutional decoder that directly outputs a $224\times 224$ solution grid. It is trained solely on supervised data and has 62M trainable parameters. (ii) Poseidon [19], a pretrained neural operator with 158M parameters. We use the Poseidon-B checkpoint and adapt it by changing the embedding and lead-time-conditioned layer normalization layers’ dimensionality to match the size of our parametrization. Similarly to the UNet, Poseidon is fine-tuned only on supervised data. (iii) PINO [29], a Fourier neural operator [28] with 33M parameters. We adapt it to accept 5-channel grid inputs and condition on the PDE operator using FiLM layers. It is trained using the same hybrid supervision and curriculum as HyPINO, including physics-informed losses.